OpenStudy (anonymous):

8xy2 - 14y -----> 2y Divide

OpenStudy (mathstudent55):

You mean this? \( 8xy^2 - 14y \div 2y = \dfrac{8xy^2 - 14y}{2y} =\)

OpenStudy (anonymous):

mhmm

OpenStudy (mathstudent55):

I am just asking if this is the problem? \( \dfrac{8xy^2 - 14y}{2y} = \)

OpenStudy (mathstudent55):

If so, then you need to divide each term of the numerator by the denominator: \( \dfrac{8xy^2 - 14y}{2y} \) \( = \dfrac{8xy^2 }{2y} - \dfrac{ 14y}{2y} = \) \( = \dfrac{2 \times 4xyy }{2y} - \dfrac{ 2 \times 7y}{2y} \) \( = \dfrac{\color{red}{\cancel{2}} \times 4x\color{green}{\cancel{y}}y }{\color{red}{\cancel{2}} \color{green}{\cancel{y}}} - \dfrac{ \color{blue}{\cancel 2} \times 7 \color{orange}{\cancel y}}{\color{blue}{\cancel{2}} \color{orange}{\cancel{y}} }\) \(=4xy - 7 \)